# Memento-II: Learning by Stateful Reflective Memory

### v1

@@ Abstract
We present a new theoretical framework for continual and experiential learning in Large Language Model (LLM) agents that integrates episodic memory with reinforcement learning. We identify reflection as the central mechanism enabling agents to adapt through interaction without back-propagation or model fine-tuning, thereby breaking the conventional separation between training and deployment. Although our prior work on Memento [zhou2025memento] and case-based reasoning LLMs [guo2025optimizing, guo2024ds] demonstrated strong empirical performance in open-ended, long-horizon tasks, the underlying learning and control principles remain theoretically unexplained. This paper further addresses this gap by introducing the Stateful Reflective Decision Process (SRDP), a formal abstraction of reflective memory dynamics, where an agent maintains episodic memory and performs two key operations: writing, which stores interaction outcomes and corresponds to policy evaluation, and reading, which retrieves relevant cases to make informed reflective decisions and corresponds to policy improvement. This equivalence allows us to analyse reflection using classical tools from dynamic programming and reinforcement learning, thereby transforming reflective memory from an empirical design pattern into a well-defined control-theoretic object. We further instantiate this stateful reflective learning by integrating retrieval with entropy-regularised policy iteration. The resulting algorithm performs soft policy iteration with guaranteed convergence. We also show that as memory expands and covers the state space more densely, the composite policy converges to the optimal solution. This theoretical framework unifies heuristic approaches such as case-based reasoning, retrieval-augmented generation with principled reinforcement learning, and provides a rigorous mathematical basis for building reflective, memory-augmented agents capable of continual adaptation and general-purpose learning.
@@ 1 Introduction
In contrast, humans exploit semantic representations, episodic memory, and reflective reuse of experience to dramatically reduce the number of examples required for effective adaptation [GershmanDaw2017EpisodicRL]. This gap suggests that learning mechanisms operating beyond purely statistical optimisation, and instead leveraging stateful memory and semantic reflection, may offer a principled route to lower-complexity learning without reliance on large-scale gradient updates.
While current LLMs remain data-hungry and inefficient in training, they nonetheless exhibit emerging cognitive abilities that enable iterative generalisation to unseen tasks with only a few examples. These abilities can be broadly categorised into three forms of what this paper called generalised reflection (Figure 2):
@@
Although these mechanisms vary in complexity from human-guided examples (Figure 2 (a)) to environment-driven feedback (Figure 2 (b)) to purely internal reasoning (Figure 2 (c)), they share a common principle: each relies on intermediate reflection to enable adaptation and generalisation beyond the training data. Such generalised reflection has been observed not only in pre-trained large language models (LLMs) [brown2020language, guo2025deepseek], but also in multimodal LLMs [driess2023palm] and non-language transformers [hollmann2025accurate].
@@
Figure 2: Three modes of generalised reflection in LLM agents: (a) in-context learning where few-shot exemplars in the prompt steer behaviour on new tasks; (b) post-output reflection via environmental feedback; (c) internal chain-of-thought and planning before emitting an answer.
In our earlier empirical work [guo2024ds, guo2025optimizing, zhou2025memento], we investigated whether reflection through memory could enable learning without parameter updates. We first introduced this paradigm with case-based reasoning LLM agents [guo2024ds], where adaptation arises from updating memory rather than fine-tuning model weights. Building on this foundation, Memento [zhou2025memento] formalised the idea within a Memory-Augmented Markov Decision Process (M-MDP) framework. Across domains such as software testing [guo2025optimizing], automated data science [guo2024ds], and deep research [zhou2025memento], this memory-driven reflection paradigm has consistently shown that LLMs can achieve continual self-improvement through accumulated experience rather than model retraining.
@@
The paper is organised as follows. We first introduce the stateful reflective decision process framework. We then present theoretical results on convergence and analyse the model’s optimality and its properties as a continuously-learnable agent. Related work is reviewed next, and the paper concludes with directions for future research.
@@ 2 Stateful Reflection as Generalised Learning Methods
Recent advances in reflective prompting, such as Tree of Thoughts [yao2023tree] and Graph of Thoughts [besta2024graph], have demonstrated the potential of structured reflection to enhance LLM reasoning. In contrast, we explore a new learning paradigm that enables models to learn from experience through reflection grounded in episodic memory, rather than through parameter fine-tuning [guo2024ds, guo2025optimizing, zhou2025memento]. To support this paradigm, we introduce the Stateful Reflective Decision Process (SRDP), which formalises continual learning as a sequence of reflective interactions with memory. We begin by introducing the core concepts, leaving the detailed mathematical formulation to the sections that follow.
Covering the three modes of generalised reflection introduced earlier (Fig. 2), we define reflection in this work as follows.
@@
Figure 1 illustrates the overall workflow of an SRDP. Episodic memory refers to the structured storage of past experiences, each recording the relationship among states, actions, and outcomes, in line with the classical notion of episodic memory in cognitive psychology [tulving1983elements]. Given a task or query represented as a state (1), the agent first enters an internal reflection stage (2), during which the LLM retrieves relevant experiences from episodic memory to guide reasoning and decision-making. Based on this reflective conditioning, the LLM produces an answer or action (3), which interacts with the external environment (4) and yields evaluative feedback, such as rewards, correctness signals, or task-level assessments. This feedback is then used to update episodic memory (5), storing new experiences that capture the consequences of the agent’s decisions. The process alternates between reading from memory for policy improvement and writing to memory for policy evaluation, enabling continual learning without modifying the underlying LLM parameters.
@@
$\pi_{t+1}=\mathcal{T}(\pi_{t})=\text{Read}\bigl(\text{Write}(\pi_{t})\bigr).$
The read step corresponds to policy improvement: the agent conditions on retrieved episodic memories, contextual exemplars, or intermediate reasoning traces to construct an improved action distribution for the current state. In LLM agents, such conditioning reshapes the effective policy directly, without parameter updates,
$\pi_{t+1}\leftarrow\mathcal{R}(\pi_{t}\mid s_{t},M_{t}),$
where $s_{t}$ denotes the current environment state and $M_{t}$ the episodic memory state. The write step corresponds to policy evaluation: the agent generates an action, plan, or answer under $\pi_{t+1}$, producing trajectories
$\tau_{t}\sim\pi_{t+1},$
@@
which is subsequently stored in episodic memory for future reads.
As we shall show in the following sections, this alternating read–write mechanism is formally analogous to policy iteration in reinforcement learning. Policy improvement is realised through memory-conditioned inference, while policy evaluation is realised through interaction-driven feedback and experience accumulation. Under mild assumptions that reading yields non-decreasing policy quality with respect to accumulated feedback and that writing provides informative evaluation signals, the iterative reflection process converges to a fixed point corresponding to an optimal or self-consistent policy.
The SRDP generalises and unifies several existing paradigms. Compared with case-based reasoning (CBR), which retrieves and adapts discrete past cases for problem solving [guo2024ds, AamodtPlaza1994], SRDP embeds such retrieval in a sequential decision process, explicitly modelling temporal dependencies and memory dynamics under the Markov assumption. In contrast to retrieval-augmented generation (RAG), which conditions generation on static external documents [Borgeaud2022RETRO], SRDP treats memory as a continually evolving component, updated through interaction and grounded in feedback. From the perspective of memory-based reinforcement learning (RL), SRDP can be viewed as replacing parameter updates with LLM reflective memory updates, allowing the policy to improve through structured recall rather than gradient-based optimisation [ramani2019short].
@@
3 Learning by Stateful Reflective Experience
@@ 3.1 Preliminary: Markov Decision Process
We begin by introducing the Markov Decision Process (MDP) [Puterman1994], the standard framework for modelling decision-making in agents. An LLM agent, viewed as a large language model capable of reasoning and acting through interaction with its environment [christianos2023pangu, yang2025agentic, wang2025tutorial, wan2024alphazero], can be naturally described within this framework. Formally, an MDP is defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$. Here, $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $\mathcal{P}(s^{\prime}\mid s,a)$ denotes the transition kernel, specifying the probability of reaching state $s^{\prime}$ after taking action $a$ in state $s$; $\mathcal{R}(s,a)\in\mathbb{R}$ represents the expected immediate reward; and $\gamma\in[0,1)$ is the discount factor. Here, the state $s$ corresponds to the current prompt or task context. Specifically, in the context of LLM reasoning, states correspond to intermediate reasoning steps [wang2025tutorial, wang2024openr], and when interacting with the external environment, they may take the form of textual descriptions of the environment state [christianos2023pangu]. The action $a$ is the model’s generated output or decision, and the transition describes how the environment or subsequent prompt evolves after the model’s response [wang2025tutorial]. A (stochastic) policy $\pi(a\mid s)$ defines a probability distribution over actions conditioned on the current state and, in turn, induces the value function that measures the expected return when following the policy from a given state.
$V^{\pi}(s)=\mathbb{E}_{\pi,\mathcal{P}}\!\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})\,\bigg|\,s_{0}=s\right],$
@@
$Q^{\pi}(s,a)=\mathcal{R}(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot\mid s,a)}\!\left[V^{\pi}(s^{\prime})\right].$
@@
$V^{\star}(s)=\max_{\pi}V^{\pi}(s),\qquad Q^{\star}(s,a)=\mathcal{R}(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot\mid s,a)}\!\left[\max_{a^{\prime}}Q^{\star}(s^{\prime},a^{\prime})\right].$
@@
$\pi_{\mathrm{LLM}}(a\mid s)\equiv p_{\mathrm{LLM}}(a\mid s;\,\theta),$
@@ 3.2 Stateful Reflective Decision Process
Building on the standard Markov Decision Process (MDP) framework introduced in the previous section, we now present the proposed Stateful Reflective Decision Process (SRDP). The classical MDP $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$ captures agent–environment interaction in its simplest form but lacks mechanisms for reflection and memory. To enable reflective, memory-driven behaviour, we extend this formulation by augmenting the agent with an episodic memory space $\mathcal{M}$, a memory retrieval policy $\mu$, and a stochastic LLM kernel $p_{\mathrm{LLM}}$ that generates context-dependent actions conditioned on both the current environment state and retrieved memory, yielding the Stateful Reflective Decision Process (SRDP). Finally, by absorbing the retrieval and LLM-generation mechanisms into effective transition and reward functions defined over the augmented state space $\mathcal{X}=\mathcal{S}\times\mathcal{M}$, we obtain a simplified induced process, referred to as the Reflected MDP. This induced process captures the full reflective dynamics of the LLM agent in a Markovian form, where a composite policy $\pi_{M}$ governs decision-making through the joint interaction of memory access and language-based reasoning. The following illustrates the relationships:
$\underbrace{\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle}_{\text{Underlying MDP}}$ $\xrightarrow{\;ext.\;}\underbrace{\Big\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{M},\mu,p_{\mathrm{LLM}}\Big\rangle}_{\text{SRDP}}\xrightarrow{\;equiv.\;}\underbrace{\Big\langle\mathcal{X}{:=}\mathcal{S}\times\mathcal{M},\ \mathcal{C},\ \mathcal{P}^{\mathrm{LLM}},\ \mathcal{R}^{\mathrm{LLM}},\ \gamma,\ \pi_{M}\Big\rangle}_{\text{Reflected MDP}}$ (1)
@@ 3.2.1 Definition
$\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{M},p_{\mathrm{LLM}}\rangle$
@@
  • $\mathcal{M}$: space of finite episodic memories, where each memory $M\in\mathcal{M}$ is a multiset of cases c=(s,a,r)c=(s,a,r) [^2] In multi-state settings, $r$ can be replaced by an estimated value $Q$, where Bellman or TD updates are propagated across memory.;
@@
At time $t$, the agent maintains episodic memory $M_{t}=\{c_{i}\}_{i=1}^{N_{t}}$, where $N_{t}$ is the number of cases in the current memeory, and observes the current environment state $s_{t}$. The decision process unfolds in two stages:
@@
    Retrieval Stage: A case $c_{t}$ is sampled from memory according to the retrieval policy
@@
    where $\mu$ is the retrieval policy (e.g., similarity-based or learned).
@@
$\pi(a\mid s,M)=\sum_{c\in M}\mu(c\mid s,M)\,p_{\mathrm{LLM}}(a\mid s,c).$ (2)
After executing the action $a_{t}$, the environment changes to $s_{t+1}\sim\mathcal{P}(\cdot\mid s_{t},a_{t})$, yields a reward $r_{t}=\mathcal{R}(s_{t},a_{t})$, and the memory updates to $M_{t+1}=\mathsf{Write}(M_{t},s_{t},a_{t},r_{t})$.
$s_{0}$$M_{0}$$c_{0}$$a_{0}$$r_{0}$$s_{1}$$M_{1}$$c_{1}$$a_{1}$$r_{1}$$s_{t}$$M_{t}$$c_{t}$$a_{t}$$r_{t}$$x_{0}$$x_{1}$$x_{t}$Environment DynamicsRetrieval Policy $\mu(c|x)$LLM Conditioning $p_{\mathrm{LLM}}(a|s,c)$Memory Write $\mathsf{Write}(M,s,a,r)$Reward Generation
@@
A key advantage of this dual-action structure is that it defines a composite policy (Eq. 2) that unifies classical reinforcement learning [SuttonBarto2018] with LLM-based decision-making [christianos2023pangu], thereby combining the strengths of both paradigms. To illustrate, consider two limiting cases. In the first case, a naive LLM produces an action $a\in c$ directly from a retrieved case $c$, effectively treating the case as a prompt. Here, the policy reduces to a standard reinforcement learning or episodic control scheme [blundell2016model, pritzel2017neural, ramani2019short], where the retrieval function $\mu$ selects the case $c$ with the highest $Q$-value, and the corresponding action is executed. In the opposite extreme, when no relevant experience exists in memory, the agent must rely entirely on the internal knowledge of the LLM to generate actions. Between these two cases lies the general setting of the SRDP, where the agent can integrate both sources of information, the reasoning capability of the LLM and the experiential knowledge stored in memory, to produce more adaptive and effective decisions.

### v2

# Memento 2: Learning by Stateful Reflective Memory

@@ Abstract
We present a theoretical study for continual and experiential learning in Large Language Model (LLM) agents integrating episodic memory with reinforcement learning. We recognise that reflection, the ability of an LLM to examine past outcomes and adjust future actions, serves as the key cognitive mechanism for continual adaptation without LLM fine-tuning. This challenges the traditional separation between training and testing in machine learning. Our empirical studies on Memento [zhou2025memento] and case-based reasoning LLMs [guo2025optimizing, guo2024ds] indicate that episodic-experience-driven reflection enables continual adaptation in open-ended, long-horizon tasks, motivating the need for a formal characterisation of the underlying learning and control mechanisms. In this paper, we fill this gap by introducing the Stateful Reflective Decision Process (SRDP), a formal abstraction of reflective memory dynamics, where an agent maintains episodic memory and performs two key operations: writing, which stores interaction outcomes and corresponds to policy evaluation, and reading, which retrieves relevant cases to make informed reflective decisions and corresponds to policy improvement. This frames reflective memory as a control theoretic object amenable to classical reinforcement learning analysis. We instantiate this Read–Write Reflective Learning by integrating retrieval into soft policy iteration, with guaranteed convergence. We formally show that as memory expands and covers the state space more densely, the composite policy converges to the optimal solution. This theoretical framework unifies heuristic approaches such as case-based reasoning, retrieval-augmented generation with principled reinforcement learning, and provides a rigorous mathematical basis for building reflective, memory-embeded LLM agents capable of continual general-purpose learning.
@@ 1 Introduction
In contrast, humans leverage semantic representations, episodic memory, and reflective reuse of experience to adapt with far fewer examples [GershmanDaw2017EpisodicRL]. This contrast suggests that learning mechanisms beyond purely statistical optimisation, based on stateful memory and semantic reflection, may offer a principled path to lower complexity learning without large scale gradient updates.
Interestingly, despite remaining data hungry and inefficient to train, current LLMs exhibit early signs of human like cognitive abilities that support iterative generalisation to unseen tasks from only a small number of examples. These emerging abilities can be broadly categorised into three forms of what we term generalised reflection, as illustrated in Figure 2:
@@
Although these mechanisms vary in complexity from human-guided examples (Figure 2 (a)) to environment-driven feedback (Figure 2 (b)) to purely internal reasoning (Figure 2 (c)), they share a common principle: each relies on intermediate reflection such as information gathering or internal computation to enable adaptation and generalisation beyond the training data. Such generalised reflection has been observed not only in pre-trained large language models (LLMs) [brown2020language, guo2025deepseek], but also in multimodal LLMs [driess2023palm] and non-language transformers [hollmann2025accurate, yu2025self].
@@
Figure 2: Three modes of generalised reflection in LLM agents, where reflection broadly refers to information gathering or internal computation conducted before an answer is produced: (a) in-context learning where few-shot exemplars in the prompt steer behaviour on new tasks; (b) post-output reflection via environmental feedback; (c) internal chain-of-thought and planning before emitting an answer.
In our earlier empirical work [guo2024ds, guo2025optimizing, zhou2025memento], we investigated whether reflection through memory could enable learning without parameter updates. We first introduced this paradigm with case-based reasoning LLM agents [guo2024ds], where adaptation arises from updating memory rather than fine-tuning model weights. Building on this foundation, Memento [zhou2025memento] formalised the idea within a Memory-Augmented Markov Decision Process (M-MDP) framework. Across domains such as software testing [guo2025optimizing], automated data science [guo2024ds, Grosnit2024KolbBasedEL], and deep research [zhou2025memento], this memory-driven reflection paradigm has consistently shown that LLMs can achieve continual self-improvement through accumulated experience rather than model retraining.
@@
The paper is organised as follows. We first introduce the stateful reflective decision process and its Read-Write Reflective Learning. We then present theoretical results on convergence and analyse the model’s optimality and its properties as a continuously-learnable agent. Related work is reviewed next, and the paper concludes with directions for future research.
@@ 2 Stateful Reflection as Generalised Learning Methods
Recent advances in reflective prompting, such as Tree of Thoughts [yao2023tree] and Graph of Thoughts [besta2024graph], have demonstrated the potential of structured reflection to enhance LLM reasoning. From a different perspective, we explore a new learning paradigm that enables models to learn from experience through reflection grounded in episodic memory, rather than through parameter fine-tuning [guo2024ds, guo2025optimizing, zhou2025memento]. To formalise this paradigm, we introduce the Stateful Reflective Decision Process (SRDP), which models continual learning as a sequence of reflective interactions with memory.
We next present the core concepts, with the full mathematical formulation deferred to later sections. Covering the three modes of generalised reflection introduced earlier (Fig. 2), we define reflection in this work as follows.
@@
Figure 1 illustrates the overall workflow of an SRDP. We begin by formally defining episodic memory as a structured repository of past interaction episodes, consistent with the classical notion of episodic memory in cognitive psychology [tulving1983elements].
Definition 2 (Episodic Memory).
An episodic memory $M_{t}$ is a finite (and growing) collection of memory items
$M_{t}=\{m_{i}\}_{i=1}^{N_{t}},\qquad m_{i}:=(s_{i},a_{i},r_{i},s^{\prime}_{i}),$
where each item $m_{i}$ records a past interaction consisting of the environment state $s_{i}\in S$, the executed action $a_{i}\in A$, the received reward $r_{i}\in\mathbb{R}$, and the subsequent state $s^{\prime}_{i}\in S$. The space of all finite episodic memories is denoted by $\mathfrak{M}$. Episodic memory evolves over time through a write operation
${M}_{t+1}=\mathrm{Write}({M}_{t},s_{t},a_{t},r_{t},s_{t+1}).$
Given a task or query represented as a state $s$ (1), the agent enters an internal reflection stage (2), during which the LLM performs a read operation on M{M} to retrieve relevant episodic memory items or their summaries. Conditioned on this retrieved information, the LLM selects an external action or generates an answer $a$ (3), which is executed in the environment (4) and produces evaluative feedback, including rewards, correctness signals, or task level assessments. Finally, the resulting feedback is incorporated into episodic memory via a write operation (5), which stores new interaction episodes.
@@
$\pi_{t+1}=\mathcal{T}(\pi_{t})=\text{\emph{Read}}\bigl(\text{\emph{Write}}(\pi_{t})\bigr).$
The Read step corresponds to policy improvement: the agent conditions on retrieved episodic memories, contextual exemplars, or intermediate reasoning traces to construct an improved action distribution for the current state,
$\pi_{t}\leftarrow\text{\emph{Read}}(\pi_{t-1}\mid s_{t},M_{t}),$
where $s_{t}$ denotes the current environment state and $M_{t}$ the episodic memory state. In LLM agents, such conditioning reshapes the effective policy directly, without parameter updates.
The Write step corresponds to policy evaluation: the agent generates an action, plan, or answer under $\pi_{t+1}$, producing trajectories
$\tau_{t}:=m_{i}=(s_{i},a_{i},r_{i},s^{\prime}_{i})\sim\pi_{t},$
@@
which is subsequently stored in episodic memory for future reads,
$M_{t+1}\leftarrow\text{\emph{Write}}(M_{t},\tau_{t}\sim\pi_{t}).$
As we shall show in the following sections, this alternating reflective read–write reflective mechanism is formally analogous to policy iteration in reinforcement learning. Policy improvement is realised through memory-conditioned inference, while policy evaluation is realised through interaction-driven feedback and experience accumulation. Under mild assumptions that reading yields non-decreasing policy quality with respect to accumulated feedback and that writing provides informative evaluation signals, the iterative reflection process converges to a fixed point corresponding to an optimal or self-consistent policy.
SRDP generalises and unifies several existing learning paradigms. Compared with case based reasoning (CBR), which retrieves and adapts discrete past cases for problem solving [guo2024ds, AamodtPlaza1994], SRDP embeds retrieval within a sequential decision making framework. This explicitly models temporal dependencies, memory evolution, and long horizon effects under a Markovian assumption, rather than treating cases as isolated problem instances.
In contrast to retrieval augmented generation (RAG), which conditions generation on static external documents or databases [Borgeaud2022RETRO], SRDP treats memory as a dynamic and stateful component. Episodic memory is continually updated through interaction with the environment and grounded in evaluative feedback, allowing retrieval to reflect the agent’s own experience rather than fixed external knowledge. From the perspective of memory based reinforcement learning, SRDP replaces parameter level updates with reflective memory updates mediated by an LLM. Policy improvement is achieved through structured recall and reflection over episodic memory rather than gradient based optimisation of model parameters [ramani2019short]. This shift enables continual adaptation while keeping the underlying LLM fixed, positioning reflection as the primary learning mechanism.
@@
3 A Model for Reflective Process over Episodic Experience
@@ 3.1 Preliminary: Markov Decision Process
To make self-contained, we begin by reviewing the Markov Decision Process (MDP) [Puterman1994], the standard framework for modelling decision-making in agents. An LLM agent, viewed as a large language model capable of reasoning and acting through interaction with its environment [christianos2023pangu, yang2025agentic, wang2025tutorial, wan2024alphazero], can be naturally described within this framework. Formally, an MDP is defined by the tuple
$\mathcal{D}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle,$
where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, and $\mathcal{P}(s^{\prime}\mid s,a)$ denotes the transition kernel, specifying the probability of reaching state $s^{\prime}$ after taking action $a$ in state $s$; $\mathcal{R}(s,a)\in\mathbb{R}$ represents the expected immediate reward; and $\gamma\in[0,1)$ is the discount factor. Here, the state $s$ corresponds to the current prompt or task context. Specifically, in the context of LLM reasoning, states correspond to intermediate reasoning steps [wang2025tutorial, wang2024openr], and when interacting with the external environment, they may take the form of textual descriptions of the environment state [christianos2023pangu]. The action $a$ is the model’s generated output or decision, and the transition describes how the environment or subsequent prompt evolves after the model’s response [wang2025tutorial]. A (stochastic) policy $\pi(a\mid s)$ defines a probability distribution over actions conditioned on the current state and, in turn, induces the value function that measures the expected return when following the policy from a given state
$V^{\pi}_{\mathcal{D}}(s)=\mathbb{E}_{\pi,\mathcal{P}}\!\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t})\,\bigg|\,s_{0}=s\right],$
@@
$Q^{\pi}_{\mathcal{D}}(s,a)=\mathcal{R}(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot\mid s,a)}\!\left[V^{\pi}_{\mathcal{D}}(s^{\prime})\right].$
@@
$V^{\star}_{\mathcal{D}}(s)=\max_{\pi}V^{\pi}_{\mathcal{D}}(s),\qquad Q^{\star}_{\mathcal{D}}(s,a)=\mathcal{R}(s,a)+\gamma\,\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot\mid s,a)}\!\left[\max_{a^{\prime}}Q^{\star}_{\mathcal{D}}(s^{\prime},a^{\prime})\right].$
@@
$\pi_{\mathrm{LLM}}(a\mid s)\triangleq p_{\mathrm{LLM}}(a\mid s;\,\theta),$
@@ 3.2 Stateful Reflective Decision Process
Building on the standard Markov Decision Process (MDP) framework introduced in the previous section, we now present the proposed Stateful Reflective Decision Process (SRDP). The classical MDP $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle$ captures agent–environment interaction in its simplest form but lacks mechanisms for reflection and memory. To enable reflective, memory-driven behaviour, we extend this formulation by augmenting the agent with an episodic memory space $M$, a memory retrieval policy $\mu$, and a stochastic LLM kernel $p_{\mathrm{LLM}}$ that generates context-dependent actions conditioned on both the current environment state and retrieved memory, yielding the Stateful Reflective Decision Process (SRDP).
@@ 3.2.1 Definition
$\mathcal{D}_{\mathrm{SRDP}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathfrak{M},p_{\mathrm{LLM}}\rangle$
@@
  • $\mathfrak{M}$: space of finite episodic memories, where each memory $M\in\mathfrak{M}$ is a multiset of cases $c=(s,a,r,s^{\prime})$ [^2] In multi-state settings, $r$ can be replaced by an estimated value $Q$, where Bellman or TD updates are propagated across memory.;
@@
At time $t$, the agent maintains episodic memory $M_{t}=\{m_{i}\}_{i=1}^{N_{t}}$, where $N_{t}$ is the number of cases in the current memeory, and observes the current environment state $s_{t}$. The decision process unfolds in two stages:
@@
    Retrieval stage: A retrieval action $c_{t}$ is sampled from the current episodic memory according to the retrieval policy
@@
    where $\mu$ may be similarity-based or learned. The action set is memory-dependent (indexer): for each $M\in\mathfrak{M}$, $\mathcal{C}(M):=M$, and each $c\in\mathcal{C}(M)$ indexes a memory case used for retrieval.
@@
$\pi^{\mu}(a_{t}\mid s_{t},M_{t})=\sum_{c_{t}\in M_{t}}\mu(c_{t}\mid s_{t},M_{t})\,p_{\mathrm{LLM}}(a_{t}\mid s_{t},c_{t}).$ (1)
After executing the action $a_{t}$, the environment changes to $s_{t+1}\sim\mathcal{P}(\cdot\mid s_{t},a_{t})$, yields a reward $r_{t}:=\mathcal{R}(s_{t},a_{t})$, and the memory updates to $M_{t+1}=\mathsf{Write}(M_{t},s_{t},a_{t},r_{t},s^{\prime}_{t})$.
$s_{0}$$M_{0}$$c_{0}$$a_{0}$$r_{0}$$s_{1}$$M_{1}$$c_{1}$$a_{1}$$r_{1}$$s_{t}$$M_{t}$$c_{t}$$a_{t}$$r_{t}$$x_{0}$$x_{1}$$x_{t}$Environment DynamicsMemory Read Policy $\mu(c|x)$LLM Conditioning $p_{\mathrm{LLM}}(a|s,c)$Memory Write $\mathsf{Write}(M,s,a,r)$Reward Generation
@@
A key advantage of this dual-action structure is that it defines a composite policy (Eq. 1) that unifies classical reinforcement learning [SuttonBarto2018] with LLM-based decision-making [christianos2023pangu], thereby combining the strengths of both paradigms. To illustrate, consider two limiting cases. In the first case, a naive LLM produces an action $a\in c$ directly from a retrieved case $c$, effectively treating the case as a prompt. Here, the policy reduces to a standard reinforcement learning or episodic control scheme [blundell2016model, pritzel2017neural, ramani2019short], where the retrieval function $\mu$ selects the case $c$ with the highest $Q$-value, and the corresponding action is executed. In the opposite extreme, when no relevant experience exists in memory, the agent must rely entirely on the internal knowledge of the LLM to generate actions. Between these two cases lies the general setting of SRDP, where the agent can integrate both sources of information, the reasoning capability of the LLM and the experiential knowledge stored in memory, to produce more adaptive and effective decisions.